A study is conducted to determine the relationship between a driver’s age and the number of accidents he or she has over a 1-year period. The data are shown here.

Drivers Age x - 16, 24, 18, 17, 23, 27, 32

No. of accidents y - 3, 2, 5, 2, 0, 1, 1

a. Draw the scatter plot.

b. Compute the value of the correlation coefficient.

c. Test the significance of the correlation coefficient at alpha = 0.01.

d. Determine the regression line equation.

e. Predict the number of accidents of a driver who is 28.


1
Expert's answer
2021-11-19T06:40:56-0500

a.





b.

correlation coefficient:

"r=\frac{\sum(x_i-\overline{x})(y_i-\overline{y})}{\sqrt{\sum(x_i-\overline{x})^2(y_i-\overline{y})^2}}"


"\overline{x}=22.4,\overline{y}=2"

"r=-0.61"


c.

Null Hypothesis: H0: r=0

Alternate Hypothesis: Ha: r≠0


test statistic:


"t=\frac{r\sqrt{n-2}}{\sqrt{1-r^2}}=\frac{-0.61\sqrt{5}}{\sqrt{1-0.61^2}}=-1.721"


"df=n-2=5"

critical value:

"t_{crit}=4.032"


Since "|t|<t_{crit}" we accept Null Hypothesis. Correlation coefficient IS NOT significantly different from zero. There IS NOT a significant linear relationship(correlation) between x and y.


d.

equation of regression line:

"y=ax+b"


"a=\frac{\sum xy-\sum x\sum y}{n\sum x^2-(\sum x)^2}=-0.1701"


"b=\frac{\sum y\sum x^2-\sum x\sum xy}{n\sum x^2-(\sum x)^2}=5.816"


"y=5.816-0.1701x"


e.

number of accidents of a driver who is 28:

"y(28)=5.816-0.1701\cdot28=1.05\approx 1" accident